Brighton, Inc., manufactures kitchen tiles. The company recently expanded, and the controller believes that it will need to borrow cash to continue operations. It began negotiating for a one-month bank loan of $600,000 starting May 1. The bank would charge interest at the rate of 1.25 percent per month and require the company to repay interest and principal on May 31. In considering the loan, the bank requested a projected income statement and cash budget for May.

 

The following information is available:

 

  • The company budgeted sales at 650,000 units per month in April, June, and July and at 500,000 units in May. The selling price is $4 per unit.

  • The inventory of finished goods on April 1 was 162,500 units. The finished goods inventory at the end of each month equals 25 percent of sales anticipated for the following month. There is no work in process.

  • The inventory of raw materials on April 1 was 61,250 pounds. At the end of each month, the raw materials inventory equals no less than 40 percent of production requirements for the following month. The company purchases materials in quantities of 67,500 pounds per shipment.

  • Selling expenses are 10 percent of gross sales. Administrative expenses, which include depreciation of $2,500 per month on office furniture and fixtures, total $160,000 per month.

  • The manufacturing budget for tiles, based on normal production of 500,000 units per month, follows:

Materials (0.25 pound per tile, 125,000 pounds, $4 per pound) $500,000

Labor                                                                                            410,000

Variable overhead                                                                         190,000 Fixed overhead (includes depreciation of $190,000)                    410,000 Total                                                                                           $1,510,000


Required:

a-1. Prepare schedules computing inventory budgets by months for production in units for April, May, and June.
a-2. Prepare schedules computing inventory budgets by months for raw materials purchases in pounds for April and May.

b. Prepare a projected income statement for May. Cost of goods sold should equal the variable manufacturing cost per unit times the number of units sold plus the total fixed manufacturing cost budgeted for the period. When calculating net sales assume cash discounts of 1 percent and bad debt expense of 0.50 percent.
Answer #1

RA) Prepare schedules computing inventory budgets by months for production in units for April, May, and June.




                                            BRIGHTON, INC.

                         Schedule Computing Production Budget (Units)

                                     For April, May, and June

                                                                  April               May           June

Budgeted sales                                      650,000        500,000       650,000

Inventory required at end of month        125,000        162,500       162,500

Total needs                                             775,000        662,500       812,500

Less: Inventory on hand at 

beginning of month                                 162,500        125,000       162,500

Budgeted production − Units                   612,500        537,500      650,000

April 500,000* 0.25= 125,000

May and June 650,000* 0.25= 162,500


RA2) 

Schedule Computing Raw Materials Inventory
Purchase Budget (Pounds)
For April and May

AprilMay
153,125134,375
53,75065,000
Total pound needs206,875199,375
61,250110,625
Balance required to purchase145,62588,750
Budgeted purchases − Pounds202,500135,000

Budgeted production − Pounds (0.25 lb. per Unit)

April: 612,500 × 0.25 = 153,125

May: 537,500 × 0.25 = 134,375

Inventory required at end of month:

April: 537,500 × 0.40 × 0.25 = 53,750

May: 650,000 × 0.40 × 0.25 = 65,000

Inventory on hand at beginning of month:

May: 110,625 = 61,250 + 202,500 – 153,125


RB)                                           BRIGHTON, INC. 

                                      Projected Income Statement

                                   For the Month of May

Sales revenue                                                             $2,000,000 

Cash discounts on sales               20,000

Estimated bad debts                     10,000

                                                                                        30,000

 Net Sales                                                                      1,970,000

Cost of Sales:

Variable cost                              1,100,000

Fixed Cost                                    410,000

                                                                                       1,510,000

Gross profit on sales                                                         460,000

Expenses:

Selling expense                             200,000

Administrative expense                  160,000

Interest expense                              7,500

                                                                                           367,500

Operating profit                                                                    92,500


2,000,000* 1%= 20,000          20,000* 0.50= 10,000

600,000+ 500,000= 1,100,000          2,000,000* 10%= 200,000


Sales revenue (500,000 Units at $4)) = $2,000,000.

Estimated bad debts (0.50 percent of gross sales) = $10,000.

Variable cost per unit = ($1,100,000 / 500,000 ) × 500,000 Units = $1,100,000.

Selling expenses (10 percent of gross sales) = $200,000.

Administrative expenses ($160,000 per month) = $160,000.

Interest expense (0.0125 × $600,000) = $7,500. 


> Accidentally put thumbs down. It's definitely a thumbs up!

KIMA Sun, Nov 28, 2021 3:27 PM
Answer #2

for RA2) 

67500* 2= 135,000

67,500* 3= 202, 500
Answer #3

For RB) fixed cost 410,00 is from fixed overhead

also for variable cost 190,000+140,000= 600,000

600,000+500,000= 1,100,000